OpenStudy (anonymous):

Evaluate 2x^2 when x=6?

OpenStudy (anonymous):

Put x = 6 \[2(6^2) = 2 \times 36\] Multiply and get the answer...

OpenStudy (anonymous):

is it 2 x^2 or (2x)^2?.. if its 2 x^2, then ans is 2* 6^2 = 2*36.. =72